The sides of a triangle are in the ratio 12:13:14. The perimeter is 156 cm. Find the lengths of the sides of the triangle.
Answer by jim_thompson5910(35256)

12x + 13x + 14x = 156

39x = 156

x = 156/39

x = 4

12x = 12*4 = 48

13x = 13*4 = 52

14x = 14*4 = 56

The three sides of the triangle are: 48 cm, 52 cm, 56 cm